The no. of roots of the equation x-2/x-2=1-2/x-1 is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Pallavi

x- 2/(x-2)=1-2/(x-1)

simplyfie it

f(x)= x3 -4x2 +5x -4=0

now f'(x) = 3x2 -8x +5

      discrimnent of this is negative  so 

  f'(x) >0 for all x

it means f(x) is an incresing function . so it will cut the x axix only at one point so number of real root of this  equation is one

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin